# DamageCAT: A Deep Learning Transformer Framework for Typology-Based Post-Disaster Building Damage Categorization

### v1

@@ 3.1 Dataset
Damage description Number of buildings Percentage Weight in loss function (Square root of inverse frequency)
Partial roof damage 7030 86.69% 1.07
Total roof damage 958 11.81% 2.91
Partial structural collapse 83 1.02% 9.9
Total structural collapse 38 0.47% 14.61
Table 1: Distribution of building damage categories with corresponding frequency weights in loss function
@@ 3.2 Model Architecture
Figure 4 presents the DAHiTrA model we are using in this work, which is proposed by Kaur et al. [26]. This hierarchical U-Net architecture is designed for satellite imagery change detection in disaster scenarios. The network processes pre-disaster and post-disaster images ($3\times 512\times 512$) through parallel ResNet-based down-sampling paths that progressively extract hierarchical features at decreasing resolutions. Each green block represents a down-sampling stage that halves spatial dimensions while increasing feature depth, creating multi-scale feature representations from $64\times 256\times 256$ down to $256\times 32\times 32$. At three intermediate levels, the difference blocks compute and analyze changes between corresponding feature maps from both temporal states. These difference blocks capture change information at multiple scales, from fine details to broader contextual changes. The decoders (purple blocks) include a dual convolutional block with intermediate normalization and 2×2$\times$2 × up-sampling operation to gradually restore spatial resolution while integrating change information from multiple scales. Skip connections (dashed lines) provide direct pathways for high-resolution spatial information to flow into the decoder. The network ends with a concatenation operation that merges the highest-level features with the original resolution features. This creates a final $5\times 512\times 512$ output that classifies each pixel into five distinct change categories (one of which is the background). This architecture effectively balances fine-grained spatial detail with hierarchical feature representation, making it suitable for identifying structural changes in post-disaster environments.
@@ 3.3 Experimental Setup
All experiments were conducted on a workstation equipped with an NVIDIA RTX 6000 GPU with 48GB of VRAM. We trained our model using the following hyperparameters: learning rate of 0.001 with a linear decay policy, input image size of 512 $\times$ 512 pixels, batch size of 8, and a maximum of 300 epochs. The Adam optimizer was used with default $\beta$ values ($\beta_{1}=0.9$, $\beta_{2}=0.999$) and $\epsilon=10^{-8}$. Early stopping was implemented with a patience parameter of 20 epochs to prevent overfitting, and the model weights from the epoch with the lowest validation loss were selected for evaluation. We implemented a custom data augmentation pipeline to increase the effective size of the dataset. Images were processed at 512 $\times$ 512 pixels and normalized with mean and standard deviation values of [0.5, 0.5, 0.5]. Our augmentation strategy included random horizontal and vertical flipping, 90/180/270 rotations, and Gaussian blur with random radius. During training, we employed a patch-based cropping approach where random crops were extracted from training images, while validation used predefined patch positions. To address the class imbalance issue, we implemented an upsampling technique that duplicated training samples containing specific class labels (particularly those with positive values in classes 2-4), with additional duplication for samples containing classes 2 or 3, effectively increasing the representation of less frequent classes in the training dataset. The model was implemented using PyTorch 2.5.1.
@@ 3.4 Quantitative Evaluation
Figure 5: Validation metrics during the training procedure.
We evaluated the model’s performance using a variety of metrics, including Intersection over Union (IoU), F1 score, precision, and recall. Table 2 summarizes the performance metrics for each damage category, as well as the overall performance across all categories. The model was trained for the full 300 epochs as specified in our experimental setup. We monitored several key metrics throughout the training process, including training loss, validation loss, IoU (Intersection over Union), and F1-score. Figure 5 illustrates the progression of various performance metrics throughout the training process over 300 epochs. The overall metrics (Figure 5 (a)) demonstrate a consistent upward trajectory, with Mean F1 score showing more rapid initial improvement compared to Mean IoU, ultimately converging to approximately 0.883 and 0.792, respectively. The training and validation loss curves (Figure 5 (c)) show typical decreasing patterns, stabilizing after 150 epochs. The minimal gap between curves indicates good generalization without overfitting.
Class-specific performance metrics reveal notable insights into the model’s ability to distinguish among the four damage categories. In particular, the partial structural collapse class attained the highest overall performance (IoU $\approx 0.839$, F1 $\approx 0.876$ at epoch 298), albeit with significant fluctuations between epochs 25 and 100. This strong yet variable outcome may stem from the visually distinctive cues associated with partial structural failure—such as exposed interior walls or partial debris fields—making it easier for the model to recognize, even if the early training stages faced instability. Meanwhile, both roof damage categories (partial and total) produced comparable IoU scores (ranging from 0.74 to 0.82) and F1 scores (0.85 to 0.90), despite having the largest number of annotated instances. This paradox suggests that visual similarity—for example, two roofs both missing shingles—can pose a classification challenge, underscoring how top-down satellite views often compress details. Moreover, the total structural collapse class underperformed slightly compared to its partial counterpart, largely due to data scarcity (only 38 labeled examples out of 8,109). Although our weighted loss and data augmentation strategies helped mitigate some imbalance issues, these results highlight the need for additional training samples to better capture the rare yet catastrophic nature of total collapses. Another key observation is that class-specific disparities were most pronounced before epoch 75, pointing to the longer learning curve for certain damage patterns, which require extensive iterative updates to achieve stable recognition. Nonetheless, all categories exceeded typical expectations for imbalanced datasets, achieving strong F1 scores (0.8519–0.9097). This robust performance across a range of damage types reinforces the efficacy of our modeling choices, including the difference blocks, transformer-based architecture, and balanced loss function, for capturing both the subtle and severe manifestations of disaster-induced building damage.
The class-specific disparities were most pronounced before epoch 75, suggesting that certain damage patterns required more training iterations to be accurately recognized. Despite implementing various techniques to address the class imbalance issue, acquiring additional samples from the underrepresented class remains necessary. Overall, contrary to typical expectations in imbalanced datasets, all categories demonstrate strong performance with F1 scores ranging from 0.8519 to 0.9097.
Category IoU F1 Precision Recall
Partial roof damage 0.7420 0.8519 0.8759 0.8292
Total roof damage 0.8128 0.8967 0.9357 0.8609
Partial structural collapse 0.8344 0.9097 0.9332 0.8874
Total structural collapse 0.7790 0.8758 0.9489 0.8132
Overall 0.7921 0.8835 0.9234 0.8477
Table 2: Performance metrics for each damage category
@@ 3.5 Qualitative Analysis
This qualitative examination reveals a high degree of concordance between predicted masks and annotated references, underlining the model’s accuracy in distinguishing multiple damage categories. Even in dense urban districts—where numerous damaged structures can sit in close proximity—the model effectively delineates individual building boundaries and assigns accurate damage labels, demonstrating fine-grained spatial discrimination. Notably, it consistently recognizes differences between partial roof damage and partial structural collapse in adjacent buildings, reflecting a nuanced grasp of visual cues associated with each damage type.
Overall, these visual findings underscore the model’s practical utility in emergency response contexts, where rapid, detailed mapping of post-disaster impacts is essential. Unlike binary or severity-only classification methods, this approach delivers categorical insights—pinpointing whether a roof is simply compromised or an entire structure is at risk of collapse. Such richer information can prove invaluable in guiding targeted interventions and resource allocation. Moreover, the consistency of these results across varied scenarios validates the robustness of our typology-based system, especially given the novelty of defining damages with such granular specificity and the absence of direct benchmarks in the current literature.
@@ 4.1 Interpretation of results
Our experimental findings underscore the viability of a categorical, deep-learning-driven approach to building damage assessment, leveraging pre- and post-disaster satellite imagery. The performance metrics for the four damage categories provide important insights into the method’s strengths and remaining challenges. Notably, the rare “partial structural collapse” class achieved unexpectedly strong results (IoU $\approx 0.8344$, F1 $\approx 0.9097$), surpassing metrics for the more common “partial roof damage” category (IoU $\approx 0.7420$, F1 $\approx 0.8519$). This discrepancy highlights that visual distinctiveness (e.g., exposed interior walls or collapsed debris fields) can override class imbalance constraints, making certain damage patterns more readily identifiable from a top-down viewpoint.
However, persistent performance gaps across categories—even after applying weighted losses and data augmentation—emphasize the intrinsic difficulties in using vertical satellite imagery for detailed damage classification. Roof damage, in particular, presents subtle visual cues (e.g., missing shingles or partial roof tarps) that can be masked by factors like shadowing, overhanging vegetation, or variations in roof color. Consequently, the two roof damage classes exhibit relatively lower IoU and F1 scores compared to structural collapse classes, whose larger-scale changes (e.g., caved-in walls, rubble) are more conspicuous in vertical imagery. Even though the 0.4,m/pixel resolution images provided valuable detail, fine-grained features remain challenging to discern due to atmospheric interference, illumination conditions, and satellite angle variations—an issue corroborated by Chen et al. [8] and Liu et al. [9].
@@ 4.2 Significance and implications
The DamageCAT framework constitutes an important advancement by elevating building damage assessment from coarse severity ratings (e.g., “minor,” “moderate,” “major”) to typology-based classifications. This novel paradigm bridges an important research gap identified by Melamed et al. [12], enabling the deep learning community to transition from generic damage labels to specific, well-defined categories (e.g., partial roof damage vs. total structural collapse). From a methodological standpoint, our transformer-enhanced U-Net architecture effectively leverages pre- and post-disaster satellite imagery to isolate disaster-induced changes, filtering out pre-existing conditions as advocated by Zheng et al. [14] and Ahn et al. [19]. This change-focused design, coupled with multi-scale difference blocks, advances the state-of-the-art by robustly detecting subtle structural variations in highly imbalanced data settings.
Equally important are the practical implications of this typology-based approach for emergency response and recovery planning. By pinpointing whether a building has partial roof damage versus a complete structural collapse, DamageCAT furnishes actionable intelligence that drastically improves resource allocation—for instance, dispatching specialized roofing teams for moderate impairments while prioritizing heavy machinery and rescue personnel for total collapses. This level of granularity facilitates targeted interventions, reduces operational uncertainties, and potentially accelerates recovery timelines compared to generic severity ratings that merely denote “light,” “medium,” or “heavy” damage. Furthermore, the ability to differentiate among multiple damage modes enhances risk assessments, enabling insurers, city planners, and engineers to prioritize critical repairs and mitigate secondary hazards (e.g., water intrusion or imminent collapse). Such data-driven insights also lend greater transparency and consistency in post-disaster decision-making, improving situational awareness for diverse stakeholders ranging from first responders to policy makers.

### v2

@@ 3.1 Dataset
Table 1 details the specific annotation criteria used to define the four building damage categories within the BD-TypoSAT dataset. For each class—Partial Roof Damage, Total Roof Damage, Partial Structural Collapse, and Total Structural Collapse—the table provides both a structural definition describing the physical state of the building and a list of corresponding satellite cues. These cues are the visual indicators observable from top-down satellite imagery that annotators used to assign a building to a specific damage category. This systematic approach ensures consistency in labeling across the dataset and provides a clear framework for understanding the distinctions between the typological damage classes.
Table 1: Annotation criteria for BD-TypoSAT damage classes
Class Structural definition Satellite cues
Partial roof damage $\leq{}50\,\%$ roof covering removed; walls intact Irregular dark patches; rafters visible; footprint unchanged
Total roof damage $50\,\%$ roof missing or complete deck loss; walls intact Entire interior exposed; clean rectangular hole; debris nearby
Partial structural collapse $\leq{}50\,\%$ load-bearing walls failed One corner flattened; L-shaped debris; shadows from remaining walls
Total structural collapse $50\,\%$ structural loss Rubble mound; indistinct footprint
@@
Table 2: Distribution of building damage categories and corresponding frequency weights.
Damage category Count Share (%) Weight ($\sqrt{1/f}$)
Partial roof damage $7030$ $86.69$ $1.07$
Total roof damage $958$ $11.81$ $2.91$
Partial structural collapse $83$ $1.02$ $9.90$
Total structural collapse $38$ $0.47$ $14.61$
@@ 3.2 Model Architecture
Figure 4 presents the DAHiTrA model we are using in this work, which is proposed by Kaur et al. [31]. This hierarchical U-Net architecture is designed for satellite imagery change detection in disaster scenarios. The network processes pre-disaster and post-disaster images ($3\times 512\times 512$) through parallel ResNet-based down-sampling paths that progressively extract hierarchical features at decreasing resolutions. Each green block represents a down-sampling stage that halves spatial dimensions while increasing feature depth, creating multi-scale feature representations from $64\times 256\times 256$ down to $256\times 32\times 32$. At three intermediate levels, the difference blocks compute and analyze changes between corresponding feature maps from both temporal states. These difference blocks capture change information at multiple scales, from fine details to broader contextual changes. The decoders (purple blocks) include a dual convolutional block with intermediate normalization and $2\times$ up-sampling operation to gradually restore spatial resolution while integrating change information from multiple scales. Skip connections (dashed lines) provide direct pathways for high-resolution spatial information to flow into the decoder. The network ends with a concatenation operation that merges the highest-level features with the original resolution features. This creates a final $5\times 512\times 512$ output that classifies each pixel into five distinct change categories (one of which is the background). This architecture effectively balances fine-grained spatial detail with hierarchical feature representation, making it suitable for identifying structural changes in post-disaster environments.
@@ 3.3 Experimental Setup
All experiments were conducted on a workstation equipped with an NVIDIA RTX 6000 GPU with 48GB of VRAM. We trained our model using the following hyperparameters: learning rate of 0.001 with a linear decay policy, input image size of 512 $\times$ 512 pixels, batch size of 8, and a maximum of 300 epochs. The Adam optimizer was used with default $\beta$ values ($\beta_{1}=0.9$, $\beta_{2}=0.999$) and $\epsilon=10^{-8}$. Early stopping was implemented with a patience parameter of 20 epochs to prevent overfitting, and the model weights from the epoch with the lowest validation loss were selected for evaluation. We implemented a custom data augmentation pipeline to increase the effective size of the dataset. Images were processed at 512 $\times$ 512 pixels and normalized with mean and standard deviation values of [0.5, 0.5, 0.5]. The dataset was split into training and validation sets with an 90:10 ratio to ensure robust model evaluation while maintaining sufficient training examples. Prior to train-validation split, a test set of 10% (213 datapoints) of the entire dataset was held-out for both qualitative and quantitative evaluation. The training was conducted four times, with different random seed set when splitting train-validaion-test sets to ensure robustness. Our augmentation strategy included random horizontal and vertical flipping, 90/180/270 rotations, and Gaussian blur with random radius. During training, we employed a patch-based cropping approach where random crops were extracted from training images, while validation used predefined patch positions. To address the class imbalance issue, we implemented an upsampling technique that duplicated training samples containing specific class labels (particularly those with positive values in classes 2–4), with additional duplication for samples containing classes 2 or 3, effectively increasing the representation of less frequent classes in the training dataset. The model was implemented using PyTorch 2.5.1.
4 Results
@@ 4.1 Quantitative Evaluation
Table 3: Validation-set performance averaged over four random data splits ($n=4$). Values are reported as mean $\pm$ standard deviation.
Category F1 IoU Precision Recall
Partial roof damage $0.809\pm 0.008$ $0.680\pm 0.011$ $0.860\pm 0.013$ $0.785\pm 0.009$
Total roof damage $0.838\pm 0.014$ $0.722\pm 0.021$ $0.917\pm 0.010$ $0.792\pm 0.032$
Partial structural collapse $0.900\pm 0.056$ $0.822\pm 0.088$ $0.960\pm 0.012$ $0.865\pm 0.093$
Total structural collapse $0.842\pm 0.061$ $0.732\pm 0.092$ $0.966\pm 0.012$ $0.781\pm 0.108$
Overall (macro) $\mathbf{0.846\pm 0.021}$ $\mathbf{0.737\pm 0.033}$ $\mathbf{0.914\pm 0.005}$ $\mathbf{0.799\pm 0.038}$
Figure 5: Convergence and variance of validation metrics. (a) shows the mean learning curves for IoU, F1 score, precision, and recall across all runs, with the shaded region representing $\pm\sigma$. (b) shows the spread of final epoch scores for these metrics using box plots. (c) and (d) show the learning curves of mIoU and F-1 scores for the four classes across the four runs.
The model was trained for the full 300 epochs as specified in our experimental setup. We monitored several key validation metrics throughout the training process. Figure 5 illustrates the progression and distribution of various performance metrics across the four independent training runs. The complete breakdown of per-class metrics (IoU, F1, precision, recall) for each training run and their progression over epochs are presented in Figures 813 in Appendix A.
Figure 5 (a) displays the convergence of overall validation metrics (Mean IoU, Mean F1, Mean Precision, and Mean Recall), averaged across the runs. These metrics generally exhibit a consistent upward trajectory, stabilizing as training progresses, indicating effective learning. The shaded regions represent one standard deviation across runs, highlighting the variability in performance. As summarized in Table 3, the model achieves a peak overall macro F1-score of $0.846\pm 0.021$ and a macro IoU of $0.737\pm 0.033$. Figure 5 (b) shows box plots that summarize the distribution of the overall IoU, F1, Precision, and Recall scores specifically from the final training epoch of each of the four runs. This subplot illustrates the consistency and spread of the model’s overall performance at the conclusion of the fixed 300-epoch training regime.
The convergence patterns of per-class F1 scores and IoU scores are detailed in Figures 5 (c) and (d), respectively. These plots show the mean and standard deviation across runs for four key damage categories: “Partial roof damage”, “Total roof damage”, “Partial structural collapse”, and “Total structural collapse”, over the 300 epochs. It is evident that class-specific disparities in learning speed and stability are most pronounced in the earlier epochs (e.g., before epoch 75).
Referring to Table 3 for the peak performance of each class (averaged best scores from four runs): The “Partial structural collapse” class achieved the highest F1-score ($0.900\pm 0.056$) and IoU ($0.822\pm 0.088$). This strong, albeit more variable (as indicated by the standard deviation and visible in early convergence in Fig. 5c,d) outcome may be attributed to its relatively distinctive visual cues.
The “Total roof damage” category also performed well, with an F1-score of $0.838\pm 0.014$ and IoU of $0.722\pm 0.021$. “Partial roof damage” yielded an F1 of $0.809\pm 0.008$ and IoU of $0.680\pm 0.011$. The performance on these roof categories, despite their visual subtleties and higher instance counts, is robust. The “Total structural collapse” class achieved an F1-score of $0.842\pm 0.061$ and IoU of $0.732\pm 0.092$. Its performance, particularly the higher variability in recall ($0.781\pm 0.108$), might be influenced by its lower representation in the dataset (38 labeled examples). Detailed per-class, per-run bar charts illustrating the spread of these best scores are available in Appendix A.
Despite the inherent challenges of class imbalance and visual similarity in satellite imagery, the model demonstrates strong performance. The convergence trends in Figure 5, coupled with robust F1 scores across all damage categories (ranging from $0.809$ to $0.900$, as per Table 3), highlight the effectiveness of our modeling choices—including the difference blocks, transformer-based architecture, and balanced loss function. While strategies like weighted loss and data augmentation helped mitigate imbalance, the results reveal areas for potential improvement, such as acquiring additional targeted samples for less represented classes.
To contextualize our model’s performance, we compare its overall macro F1-score with those reported in several recent studies on building damage assessment using satellite imagery. Table 4 presents this comparison with other methods, many of which focus on binary or ordinal damage classification. It is important to note that direct comparisons can be nuanced due to differences in datasets, evaluation protocols, and the specific definitions of damage categories used in each study. Consequently, we do not report per-class metrics: the typology-based damage classes introduced in DamageCAT share no one-to-one correspondence with the binary or ordinal labels used in prior work, so a class-level comparison would be conceptually misleading.
Table 4: Overall F1-score comparison between our model and prior work
Method F1 Score
UNet+BDANet +morphology, Neto and Dantas [10] 0.7998
MTF, Weber and Kané [21] 0.7417
HRTBDA, Chen et al. [8] 0.8071
Bouchard et al. [5] 0.8460
Wang et al. [13] 0.8322
Liu et al. [9] 0.7017
Ours 0.8458
@@ 4.2 Qualitative Analysis
This qualitative examination shows general agreement between predicted masks and annotated references. This suggests the model can distinguish multiple damage categories. Even in dense urban areas, where many damaged structures are close together, the model appears to delineate individual building boundaries and assign damage labels with reasonable accuracy. This indicates an ability for spatial discrimination. Notably, the model consistently recognizes differences between partial roof damage and partial structural collapse in adjacent buildings. This suggests it can learn some visual cues associated with each damage type.
Overall, these visual findings demonstrate the model’s practical utility in emergency response contexts, where rapid, detailed mapping of post-disaster impacts is essential. Unlike binary or severity-only classification methods, this approach delivers categorical insights—pinpointing whether a roof is simply compromised or an entire structure is at risk of collapse. Such richer information could be valuable for guiding targeted interventions and resource allocation. Moreover, the consistency of these results across varied scenarios suggests some robustness in our typology-based system. This is important because defining damages with such specific detail is a novel aspect of our system, and direct benchmarks are not available in current literature.
Figure 7: Model transferability evaluation across multiple hurricane events. The figure shows prediction results on Hurricane Harvey (column 1), Hurricane Florence (column 3), and Hurricane Michael (columns 2, 4–7). Each row displays pre-disaster imagery, post-disaster imagery, model predictions, and overlay visualization respectively.
To evaluate the model’s transferability beyond Hurricane Ida, we tested it on satellite images from three other hurricanes: Harvey (2017), Florence (2018), and Michael (2018), as shown in Figure 7, which all come from the dataset xBD [40]. The results demonstrate the model’s ability to generalize across different hurricane events, locations, and types of buildings without needing extra training. The model successfully found damaged buildings in the Houston area from Hurricane Harvey (leftmost column) and in North Carolina’s coastal areas after Hurricane Florence (column 3). The results from Hurricane Michael (columns 2 and 4–7) also show it worked well across Florida’s Panhandle region. Although the model’s confidence sometimes varied, likely due to factors like image quality, different seasons, and local building styles, it consistently classified the type of damage correctly across all these hurricanes. This consistency suggests our framework could be a valuable tool in real-world disaster response. However, quantitative evaluation of these cross-disaster results was not feasible due to fundamental differences in labeling criteria between our typology-based damage categories and the severity-based annotations employed in the xBD dataset.
@@ 5.1 Interpretation of results
Through experimental validation, we establish that a categorical, deep-learning-driven approach to building damage assessment can effectively leverage pre- and post-disaster satellite imagery. The performance metrics for the four damage categories provide important insights into the method’s strengths and remaining challenges. Notably, the rare “partial structural collapse” class achieved unexpectedly strong results (IoU $\approx 0.8344$, F1 $\approx 0.9097$), surpassing metrics for the more common “partial roof damage” category (IoU $\approx 0.7420$, F1 $\approx 0.8519$). This discrepancy highlights that visual distinctiveness (e.g., exposed interior walls or collapsed debris fields) can override class imbalance constraints, making certain damage patterns more readily identifiable from a top-down viewpoint.
However, persistent performance gaps across categories—even after applying weighted losses and data augmentation—emphasize the intrinsic difficulties in using vertical satellite imagery for detailed damage classification. Roof damage, in particular, presents subtle visual cues (e.g., missing shingles or partial roof tarps) that can be masked by factors like shadowing, overhanging vegetation, or variations in roof color. Consequently, the two roof damage classes exhibit relatively lower IoU and F1 scores compared to structural collapse classes, whose larger-scale changes (e.g., caved-in walls, rubble) are more conspicuous in vertical imagery. Even though the 0.4 m/pixel resolution images provided valuable detail, fine-grained features remain challenging to discern due to atmospheric interference, illumination conditions, and satellite angle variations—an issue corroborated by Chen et al. [8] and Liu et al. [9].
@@
While our typology-based approach represents an advancement in damage classification granularity, it builds upon decades of prior work in remote sensing-based damage assessment. Earlier studies have attempted various forms of detailed damage categorization, though often constrained by the resolution limits of available imagery and computational capabilities of their time. Our framework’s performance, while promising, still faces some of the fundamental challenges identified in seminal works: the inherent limitations of vertical imagery for detecting certain damage types, variations in classification accuracy across different urban morphologies, and the persistent challenge of rare damage cases. These constraints underscore that while deep learning enables more nuanced damage classification, careful consideration of operational contexts and potential accuracy trade-offs remains essential for practical deployment.
@@ 5.2 Significance and implications
The DamageCAT framework advances building damage assessment by moving beyond coarse severity ratings (e.g., “minor,” “moderate,” “major”) to typology-based classifications that specify the nature and location of structural damage. This approach addresses a key research gap identified by Melamed et al. [12], enabling transition from generic damage labels to well-defined categories such as partial roof damage versus total structural collapse. Methodologically, our transformer-enhanced U-Net architecture effectively isolates disaster-induced changes by comparing pre- and post-disaster satellite imagery, filtering out pre-existing conditions as advocated by Zheng et al. [19] and Ahn et al. [24]. The multi-scale difference blocks enable robust detection of subtle structural variations even in highly imbalanced datasets.
The practical implications for emergency response are significant. Traditional severity classifications provide limited actionable information, whereas typological damage assessment enables targeted resource allocation. For example, buildings with partial roof damage require roofing specialists and materials, while those with total structural collapse need heavy machinery and search-and-rescue teams. This specificity reduces operational uncertainties and accelerates response times compared to generic severity ratings.
Beyond immediate response, typological classification enhances long-term recovery planning and risk assessment. Engineers can prioritize critical repairs based on specific damage modes, while insurers gain more precise loss estimates. City planners can identify patterns of structural vulnerability to inform future building codes and disaster preparedness strategies. The framework’s granular damage descriptions also improve transparency in post-disaster decision-making, providing clear, consistent information for stakeholders ranging from first responders to policy makers.